OpenStudy (anonymous):

test the convergence Xn= 1+1/2!+1/3!.....1/n!

OpenStudy (anonymous):

see its bounded '\(1 \le X_n \le 1+c \) :P check c as uper bound also its monotonic \(a_n > a_{n-1}\) so bounded +monotonic =convergence :D

OpenStudy (anonymous):

awhhh okay okay

OpenStudy (anonymous):

u dnt need to find c though :D

OpenStudy (anonymous):

since 1/n! goes to zero then its bounded on half closed set

OpenStudy (zarkon):

1/n also goes to zero..does that tell you anything about 1+1/2+1/3+...+1/n as n goes to infinity

OpenStudy (anonymous):

yep , its also converge

OpenStudy (zarkon):

no...it is the harmonic series...it diverges

OpenStudy (anonymous):

bounded +monotonic :O

OpenStudy (zarkon):

if the series is bounded and monotonic then it converges

OpenStudy (zarkon):

you are just bounding a term and not the whole series

OpenStudy (anonymous):

ok let me use the def :)

OpenStudy (reemii):

there's a criterion which involves \(\lim_{n\to\infty}\frac{a_{n+1}}{a_n}\)..

OpenStudy (anonymous):

A sequence X=\(\large (x_n)\) in R is said to be converge to \(x\in R\), or x said to be a limit of \(x_n\) if for every \(\epsilon >0 \) there exist a natural number\( K( \epsilon )\) such that for all n\(\ge \) \( K( \epsilon )\) the terms \(x_n \) satisfy \( |x_n-x | \) < \(\epsilon \) if a sequence has a limit , we say that the sequence is convergent ,if it has no limit , we say that the sequence is divergent

OpenStudy (anonymous):

wew ! according to this i say its conv. also 1/n conv what do u say ?

OpenStudy (zarkon):

yes @reemii that is one way to do it if one does not recognize the sum

OpenStudy (zarkon):

the original does converge...I know the sum. \[\sum_{n=1}^{\infty}\frac{1}{n}\]diverges so they way you are thinking is wrong

OpenStudy (anonymous):

why lol

OpenStudy (anonymous):

>.> hehe its series and i prove for sequence , sorry

OpenStudy (anonymous):

o_o i hv 4 more assignments to write

OpenStudy (anonymous):

in 2 hours

OpenStudy (anonymous):

so what do i write here >_<

OpenStudy (anonymous):

harmonic sequence diverges so it does not converge?

OpenStudy (anonymous):

1/n div 1/n! conv but lets reason it again , dnt write the comment above

OpenStudy (anonymous):

okay?

OpenStudy (anonymous):

.-.

OpenStudy (anonymous):

\( \large x_n =\sum \frac {1}{n!} \) its convergent , enough to show by Induction n^2 <n! for n\(\ge\) 4 but first u need to know 1/n^2 is conv. also ( u can prove it :) ) so \(\large 0<\frac {1}{n!}<0<\frac {1}{n^2}\) for n>=4 >.>

OpenStudy (anonymous):

\(\large x:=\frac {1}{n!} \) \(\large y:=\frac {1}{n^2} \) then when \(n\ge 4\) we have \(\large 0\le \frac {x_n}{y_n }=\frac {n^2 }{n!} <\frac {1}{n-2} \rightarrow 0 \)

OpenStudy (anonymous):

.-. i cant remember this fully lol but i guess its correct hehe

OpenStudy (anonymous):

then by \(\Huge \color{red}{\text { Limit Comparison Test }}\) Apllies the conv "u should know this test :D "

OpenStudy (anonymous):

done :D

OpenStudy (anonymous):

bravo ^^

OpenStudy (anonymous):

ikr ^^ thx to @Zarkon to notice that i was working for sequence :D